OpenStudy (anonymous):

Use the quadratic formula to solve x2 - 2x - 3 = 0

OpenStudy (anonymous):

\[ x^2 - 2x - 3= (x+1)(x-3) \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

D = b^2 - 4ac D = 4 - 4*(-3)*1 D = 16 Therefore sqrtD = 4 Putting in formula, = -b +/- (sqrt D)/2a = (2 (+/-) 4 )/ 2 = 1+/- 2 = 3 or -1